I am sorry I need help with
Find the zeros of the function
f(x)= x^2 - 10x + 74
I typed it incorrect last time. I am sorry. Please help me.

Answer by stanbon(75887)

Find the zeros of the function
f(x)= x^2 - 10x + 74
Use the quadratic formula:
x = [10 +- sqrt(100 -4*1*74)]/2
x = [10 +- sqrt(-196)]/2
x = [10 +- 14i]/2
x= 5 + 7i or x = 5 - 7i
==========================
Cheers,
Stan H.